7. Water runs into a fountain, filling all the pipes, at a steady rate of 0.750 m³/s. (a) How fast will it shoot out of a hole 4.50 cm in diameter? (b) At what speed will it shoot out if the



diameter of the hole is three times as large?

1
Expert's answer
2022-06-13T08:24:08-0400

"\dot V=vS=\frac{\pi vd^2}4,\implies"

"v=\frac{4\dot V}{\pi d^2}=472~\frac ms,"

"v_1=\frac v9=52~\frac ms."